OpenStudy (anonymous):

help me please c: will give medal and fan c:

OpenStudy (anonymous):

A rectangular cardboard has dimensions as shown. The length of the cardboard can be found by dividing its area by its width. What is the length of the cardboard in inches? a)9 and 41 over 51 b)10 and 1 over 4 c)36 and 5 over 6 d)174 and 29 over 48

OpenStudy (anx):

The area of a rectangle is A= l(w) You're given: 41 2/3 = l(4 1/4) turn to decimals to make easier 41.666667 = l (4.25) divide both sides by 4.25 l = 9.80392 well you know that the answer will start with 9 then just find the decimals of fractions 41/51 = .80392 boom first try Answer is A 9 41/51

OpenStudy (anonymous):

thank you so much @anx

OpenStudy (anx):

no problem.